OpenStudy (anonymous):

write an equation of a line slope 6;through (0,4)

OpenStudy (saifoo.khan):

y = 6x + 4

OpenStudy (anonymous):

You need to put the problem into point intercept form

OpenStudy (anonymous):

yeaa

OpenStudy (saifoo.khan):

you want an equation. :P

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

No I know the equation. :) It is y-y1=m(x-x1) So : y-4=6(x-0) Now you simplify: y= 6x+4 :D

OpenStudy (anonymous):

ok look these explanations suck. the form is y=mx+b m is your slope you you already know the number that goes there is 6 to find b sub in the x and y values of the point into the equation so 4=6(0)+b you find that b is 4 so now you have all of the components to make the equation that is y=6x+4